Physical Chemistry

For water at 250C, 2*10-7 moles per litre is the correct answer for which one of the following-
1.[H+] + [OH-] 2. [H+]2 3. [OH-]

Tanusri Bose
7 Years agoGrade

1 Answer

Rituraj Tiwari
5 Years ago
At25oC, For pure waterConcentration of[H+]=10−7moles per litre
Concentration of[−OH]=10−7moles per litre
aspH=pOH=7
Thus[H+]+[−OH]=10−7+10−7=2×10−7moles/L